OpenStudy (idealist10):

Find the curvature at the point (x, y) on the ellipse x^2/9+y^2/4=1.

OpenStudy (idealist10):

@myininaya

OpenStudy (idealist10):

@myininaya

myininaya (myininaya):

so have you thought of doing using parametric equations?

OpenStudy (idealist10):

Yes. Here's my work: x^2/a^2+y^2/b^2=1 x(t)=acos(t) y(t)=bsin(t) ----------------- x(t)=3cos(t) y(t)=2sin(t) in this case ---------------- k=abs(x'y"-y'x")/[(x')^2+(y')^2]^(3/2) x'(t)=-3sin(t) y'(t)=2cos(t) x"(t)=-3cos(t) y"(t)=-2sin(t) k=6/(9sin^2 t+4cos^2 t)^(3/2) And I'm stucked.

myininaya (myininaya):

\[k=\frac{6}{(9\sin^2(t)+4\cos^2(t))^\frac{3}{2}} \\ \text{ there isn't much you can do with this except } \\ k=\frac{6}{(5 \sin^2(t)+4\sin^2(t)+4 \cos^2(t))^\frac{3}{2}}=\frac{6}{(5 \sin^2(t)+4(1))^\frac{3}{2}} \\ k=\frac{6}{(5 \sin^2(t)+4)^\frac{3}{2}}\]

myininaya (myininaya):

that looks a lil bit prettier

OpenStudy (idealist10):

But that's not the answer. The answer in the book says \[\frac{ 162 }{ (81-5x^2)^{3/2} }\]

myininaya (myininaya):

oh! they want it back in terms of x and y (or I mean just x) \[k=\frac{6}{(5(\frac{y}{2})^2+4)^\frac{3}{2}} \\ =\frac{6}{(\frac{5y^2}{4}+4)^\frac{3}{2}} \\ =\frac{6}{(5(1-\frac{x^2}{9})+4)^\frac{3}{2}}\] so we are going to play with this

myininaya (myininaya):

\[=\frac{6}{(5 -5 \frac{x^2}{9}+4)^\frac{3}{2}}=\frac{6}{(9-\frac{1}{9}5x^2 )^\frac{3}{2}}\]

myininaya (myininaya):

now we have a compound fraction

myininaya (myininaya):

so we need to find a way to get rid of that we can do this... \[=\frac{6}{(\frac{1}{9} \cdot (18-5x^2))^\frac{3}{2}}\]

myininaya (myininaya):

guess what we can do by law of exponents?

myininaya (myininaya):

oops 81 sorry I'm dyslexic sometimes

myininaya (myininaya):

\[=\frac{6}{(\frac{1}{9} \cdot (81-5x^2))^\frac{3}{2}}\]

myininaya (myininaya):

you know since 1/9*81=9 and 1/9 *5=5/9

myininaya (myininaya):

but anyways by law of exponents we can say \[=\frac{6}{(\frac{1}{9} \cdot (81-5x^2))^\frac{3}{2}}=\frac{6}{(\frac{1}{9})^\frac{3}{2}(81-5x^2)^\frac{3}{2}}\]

myininaya (myininaya):

but you can simplify from here

myininaya (myininaya):

I bet*

myininaya (myininaya):

but if not let me know

OpenStudy (idealist10):

But how did you get \[\frac{ 6 }{ (5(1-\frac{ x^2 }{ 9 })+4)^{3/2} }\]

myininaya (myininaya):

\[\frac{x^2}{9}+\frac{y^2}{4}=1 \\ \frac{y^2}{4}=1-\frac{x^2}{9}\]

myininaya (myininaya):

the ellispe equation

myininaya (myininaya):

remember you called y=2 sin(t) so sin(t)=y/2

OpenStudy (idealist10):

Oh, and how did you get \[\frac{ 6 }{ (9-\frac{ 1 }{ 9 }5x^2)^{3/2} }\]

myininaya (myininaya):

\[5 \sin^2(t)+4 \\ 5(\frac{y}{2})^2+4 \\ 5 \frac{y^2}{4}+4 \\ 5(1-\frac{x^2}{9})+4 \\ 5-5 \cdot \frac{x^2}{9}+4 \\ 9-5 \frac{x^2}{9} \\ \frac{1}{9}(9^2-5 x^2)\]

OpenStudy (idealist10):

Oh, and how did you get \[\frac{ 6 }{ (9-\frac{ 1 }{ 9 }5x^2)^{3/2} }\]

OpenStudy (idealist10):

Oh, and how did you get \[\frac{ 6 }{ (9-\frac{ 1 }{ 9 }5x^2)^{3/2} }\]

myininaya (myininaya):

distribute and added like terms above

OpenStudy (idealist10):

Thank you so much for the help, @myininaya !! I'm so sorry that in the morning, my computer was slow.